0.05M glucose solution is isotonic with ______ M of NaCl

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Both should have same Osmotic Pressure.
Therefore,
0.05*ST = 2*CST
c = 0.05/2 = > 1/40 M
Hence to be isotonic it shoud have 1/40 molarity concentration.